((4/(3r-1))-4)/((4/(3r-1))+4)
1
Expert's answer
2012-12-13T08:13:02-0500
((4/(3r-1))-4)/((4/(3r-1))+4)=[ 4/(3r-1)-4(3r-1)/(3r-1)]/ [ 4/(3r-1) +4(3r-1)/(3r-1) ]= =[ (4-4(3r-1) )/(3r-1) ]/ [ (4+4*(3r-1)
)/(3r-1) ]= =[ (4-12r+4)/(3r-1) ]/ [ (4+12r-4)/(3r-1) ]= =[ (8-12r)/(3r-1)]/
[12r/(3r-1)]=(8-12r)/(12r)=8/(12r) -1=2/(3r) -1=(2-3r)/3r